OpenStudy (anonymous):

the system of equations is coincident. what are the mising valves? -2+5y= - 12 -6x + y =

OpenStudy (anonymous):

hello you need help

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

what are the missing valves? -2x + 5y =-12 -6x + ______y =____________

OpenStudy (anonymous):

ok one sec let me solve it

OpenStudy (anonymous):

are you still there

OpenStudy (anonymous):

yea sorry i was helpping another

OpenStudy (anonymous):

no sorry i dont know

OpenStudy (anonymous):

i am sorry i wish i new

OpenStudy (anonymous):

hello you need help